the supply and demand functions for a product are defined as Qs =100 and
Qd =50P^1/2 + 10
i)find the equilibrium price and quantity

ii)sketch a graph to show the results in (i)above

iii)Assume the equilibrium price is 80 pesewas (p=0.80)
find the PES and PED associated with a small increase in price from current price of 80 pesewas.
iv)What is the incidence of a tax of Ghs 0.6 imposed on the producer?
1
Expert's answer
2020-03-26T12:07:53-0400

Solution:

i) Qs=Qd


"50 \times {\sqrt{p}}+10=100"

"p=\frac {81}{25}"

"p=3.24"

"Q=100"

ii)



iii)


"PES=Qs{\prime}(p) \frac {p}{Qs(p)}=0"


"PED=Qd{\prime}(p) \frac {p}{Qd(p)} = 0.41"

iv)

"0.6\times 100=60"